# Low Variance Off-policy Evaluation with State-based Importance Sampling

### v4

@@ 7.1 Deterministic lift domain
An analysis comparing state-based estimators to their traditional counterparts (see Table 1) demonstrates a near universal benefit in the deterministic domain, with the only exception being SDRE. SIS is the best performing estimator across all domain sizes, followed by IS and SPDIS depending on the domain size. SDRE and SSDRE have particularly poor estimates because the domain does not have a stationary distribution. Analysis of the residuals (see Figure 1a) also shows that: a) state-based methods have successfully reduced the variance compared to their traditional counterparts; b) SPDIS strongly reduces the residual of PDIS for large domain sizes; and c) INCRIS and SINCRIS have large residuals, which implies that dropping time steps may not be suitable for the deterministic domain (e.g. the first time step always determines the final reward).
Due to the deterministic domain being easy to model, the doubly robust methods have MSEs close to zero in this domain so are not further analysed.
(a) Deterministic lift domain
(b) Stochastic lift domain
Figure 1: Residuals ($y$-axis), defined as $\hat{G}-\mathcal{G}$, as a function of the domain size ($x$-axis) in the lift domains. Estimates are based on 1,000 episodes. Residuals are represented by their mean $\pm$ standard error over 50 independent runs for the deterministic lift domain and over 200 independent runs for the stochastic domain.
Table 1: Mean squared error (MSE) for off-policy evaluation estimators of the expected return in the deterministic lift domain based on 50 independent runs for each domain size. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size. All SIS estimators shown use the Q-value based algorithm for negligible state set identification.
(a) 100 episodes
Domain size
IS
SIS
PDIS
SPDIS
INCRIS
SINCRIS
SDRE
SSDRE
7 0.0477 0.0324
0.3732
0.2393
3.9565
3.9489
309.0670
309.9788
9 0.1084 0.0285
1.3757
0.3131
6.3628
6.2632
300.3519
304.1609
11 0.3165 0.0338
6.1956
0.6621
9.0336
8.4515
461.0101
474.6804
13 0.5054 0.0246
14.1812
0.6062
17.2282
12.4712
412.2747
430.6147
15
0.8874
0.0273
32.1658
0.8837
29.8467
15.3794
553.0677
582.1068
17
2.0117
0.0250
99.5122
0.8980
39.8329
17.1526
558.1893
587.1945
(b) 1,000 episodes
Domain size
IS
SIS
PDIS
SPDIS
INCRIS
SINCRIS
SDRE
SSDRE
7 0.0067 0.0019
0.0398
0.0121
4.0024
4.0007
532.7567
536.3787
9 0.0153 0.0019
0.1766
0.0215
6.1964
6.1765
577.3089
584.4657
11 0.0353 0.0029
0.7087
0.0600
8.9408
8.9225
674.0772
680.1959
13
0.0691
0.0026
1.8038
0.0688
12.5475
12.4779
652.6004
660.2387
15
0.1389
0.0033
5.3474
0.1378
16.1902
16.1558
803.3450
813.4997
17
0.3026
0.0031
15.1566
0.1303
23.7084
20.1534
758.2237
764.6283
7.2 Stochastic lift domain
In addition to the deterministic lift domain, we also show a stochastic lift domain in which the evaluation policy and the state transitions are stochastic rather than deterministic. The evaluation policy $\pi_{e}$ takes the best action with probability $1-\delta$ and the other action with probability $\delta$, and the state transitions dynamics are formulated as $\mathcal{T}^{\prime}(s,a,s^{\prime})=1-\delta$ where $\mathcal{T}(s,a,s^{\prime})=1$ and $\mathcal{T}^{\prime}(s,a,s^{\prime})=\delta$ otherwise – this includes lift states and non-lift states. The stochastic dynamics imply that the agent does not necessarily proceed forwards through the lift but may at some times go from one side of the line to the other side of the line. The true expected return is now negative due to typically not being able to take a straight path towards the goal, and the larger the domain the stronger this negative effect becomes.
Table 2: Mean squared error (MSE) for off-policy evaluation estimators of the expected return in the stochastic lift domain based on 200 independent runs for each domain size. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each domain size.
(a) 100 episodes
Domain size
WIS
WSIS
WPDIS
WSPDIS
WINCRIS
WSINCRIS
SDRE
SSDRE
WDR
WDRSIS
7
11.8501
9.9437
7.9853
6.8646
1.6110 0.9388
141.2766
143.0795
6.1663
5.6135
9
10.4314
8.8458
6.3404
5.3060
0.5713 0.6468
163.2390
168.3226
4.0534
3.9546
11
11.5886
7.6150
6.5636
4.6576
0.8474 1.4792
161.3365
163.2882
3.3082
3.1341
13
10.3529
11.1001
7.8276
5.2161
2.3283
3.1879
190.9451
197.5445
2.9028 2.6901
15
10.1513
8.2299
8.6465
3.9065
6.9703
6.5948
213.4919
222.7848
2.3015 1.9605
17
10.5048
5.2746
12.9510
2.9489
12.5488
10.0513
246.1740
254.5904
1.9964 1.4650
(b) 1,000 episodes
Domain size
WIS
WSIS
WPDIS
WSPDIS
WINCRIS
WSINCRIS
SDRE
SSDRE
WDR
WDRSIS
7
10.2238
8.4832
6.8271
5.3853
1.0958 0.2618
208.7413
209.8354
5.2572
5.0844
9
9.4737
7.7326
5.3924
4.0441
0.2973 0.6929
243.1366
245.1004
2.8067
2.7486
11
9.7285
6.2044
5.7219
3.0552
0.2002
2.3050
250.3149
251.5843
2.3016 1.9838
13
8.4436
9.4967
4.4449
3.2916
0.8958
4.8960
266.7911
268.6634
1.3459 1.3354
15
18.0453
4.2831
6.2259
1.7004
2.6550
9.0162
280.6434
282.6849
1.2258 0.8797
17
8.2474
4.8129
5.2761
1.8151
4.7157
13.0056
305.7465
307.9191
0.8385 0.4820
Results for state set identification are analogous to the deterministic lift domain: WSIS (Q-based) is successful at identifying the lift states as negligible states and together with WSIS (Lift states) has a superior performance compared to WSIS (Covariance testing) and WIS (see Table 6 in Appendix C); and WSIS (Covariance testing) typically drops two lift states (the maximal cardinality given the computational budget). Table 7 demonstrates how dropping states is also beneficial for doubly robust methods and again the Q-based state identification works to identify the lift states.
An analysis comparing state-based estimators to their traditional counterparts (see Table 2) demonstrates a benefit in the stochastic domain. This time doubly robust methods are also analysed as the error of the DR estimator is now non-zero. WSIS and WSPDIS consistently outperform their traditional counterparts. For large domain sizes, WDRSIS is the best performing estimator followed by WDR and WSPDIS. It is notable that state-based methods are not adversely affected by the domain size, which is explained by all the lift states being dropped and therefore the maximal number of steps $M_{B}$ in non-negligible states remaining constant regardless of domain size. For small domain sizes, WINCRIS and WSINCRIS have the best performance, indicating the benefit of dropping time steps; this is a clear difference compared to the deterministic domain, which is because now a) trajectories are much longer, requiring more variance reduction; and b) both the length of the trajectory and the ability of the agent to switch between the different lifts reduces the importance of any given time step, making time steps more amenable to being dropped. The lower performance of WINCRIS and WSINCRIS in larger domain sizes may be attributed to the maximal number of time steps; that is, it is not computationally feasible to check all $k\in[1,H)$ so a setting of $k\in[1,10]$ was used instead. Figure 1b further demonstrates that the state-based methods have successfully reduced the variance compared to their traditional counterparts except for WINCRIS, which has already significantly reduced the variance. As in the deterministic lift domain, the poor performance of SDRE estimators in the stochastic lift domain is attributed to the domain not having a defined stationary state distribution.
7.3 Inventory management
The third domain investigated is the inventory management problem [16]. The task of the agent in the inventory management problem is to purchase items to make optimal profits selling the items. In doing so, the agent must balance supply with demand while having at most $S$ items in the inventory. The state is the current inventory while the action is the added chosen number of items to supply. States corresponding to an almost full inventory (i.e. close to $S$ items) are not pure lift states but similar to lift states the actions chosen are unimportant – that is, the inventory will be close to or equal to $S$ regardless of the action. For each item, the purchasing cost is set to $2.49$, the sale price to $3.99$, and the holding cost to $0.03$. Demand for items follows a Gaussian with mean $S/4$ and standard deviation $S/6$, and $S=10$ in the experiments. Each episode consists of 100 steps, and consistent with the rest of our experiments an undiscounted setting ($\gamma=1$) is considered. The evaluation policy, $\pi_{e}$, was trained for 500 episodes on the MDP and has an expected return of $\mathcal{G}=1471.1041$. The behaviour policy, $\pi_{b}$, is uniform random. As the domain is high variance due to the varying demand, we opt for a relatively high number of 50 experiment runs and instead of manipulating the effective horizon we analyse the improvement in performance across different sample sizes (i.e. numbers of episodes).
Table 3: Normalised mean squared error (MSE) for Monte Carlo estimators of the expected return in the inventory management domain based on 50 independent runs for each number of episodes. Normalisation is based on division by the squared range. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each number of episodes.
Episodes
WIS
WSIS
WPDIS
WSPDIS
WINCRIS
WSINCRIS
SDRE
SSDRE
WDR
WDRSIS
100
0.0490
0.0088
0.0278
0.0074
0.0641
0.0109
0.0160
0.0077
0.0173
0.0138
250
0.0651
0.0080
0.0335
0.0080
0.0546
0.0118
0.0161
0.0070
0.0132
0.0137
500
0.0652
0.0113
0.0340
0.0105
0.0456
0.0124
0.0163
0.0067 0.0098
0.0114
1000
0.0820
0.0122
0.0293
0.0119
0.0313
0.0128
0.0162
0.0066 0.0117 0.0104
An analysis comparing state-based estimators to their traditional counterparts (see Table 3) demonstrates a near universal benefit in the inventory management domain, with 2– to 8–fold improvements being observed (see Table 3), with the exception of DR which already has high accuracy. WSPDIS performs best for 100 episodes while SSDRE performs best for 250 to 1,000 episodes. The high performance of SSDRE is attributed to a clearly defined stationary distribution. The performance of SSDRE and SDRSIS depending on the sample size suggests that these state-based methods provide improved convergence and efficiency. Other high-performing algorithms are WSIS, WSINCRIS, WDR, and WDRSIS. Figure 2a demonstrates that state-based techniques reduce the variance and yield a smaller residual.
7.4 Taxi
The fourth domain is the variant of the taxi domain mentioned in Liu et al. (2018) [8], which has a much larger state space with $|\mathcal{S}|=2000$ and can have infinite horizon. The domain includes a taxi agent that has to pick up and drop off passengers who arrive at random corners in a $5\times 5$ grid based on $6$ actions, namely going a step north, west, south, east, picking up or dropping off a passenger. The taxi agent receives a reward of $20$ when picking up a passenger or dropping off a passenger at the correct location, and a reward of $-1$ otherwise. The implementation of the domain based on code of Liu et al. (2018) [8] (see {https://github.com/zt95/infinite-horizon-off-policy-estimation/tree/master/taxi}). The evaluation policy represents a policy formed after 1,000 iterations of Q-learning and the behaviour policy represents a policy formed after 950 iterations of Q-learning. The domain is of interest to investigate the scalability to large state-action spaces and large effective horizon, and it is of interest to manipulate the effective horizon in an additional domain which in contrast to the lift domains has a stationary distribution. Due to the increased computational expense, the number of runs and the number of episodes per run are limited to 20 and 100, respectively.
As in the previous domains, the results for the taxi domain (see Table 4) also show a near universal benefit for state-based estimators compared to the traditional counterparts, although the effect is not as pronounced as for example in the inventory management domain. Analysis of the effective horizon, i.e. the episode length, shows that for an effective horizon of 10, WSIS is the highest performing followed by WSINCRIS and WINCRIS while for effective horizons between 50 and 1,000, the SSDRE is the highest performer followed by SDRE and WSPDIS. The high performance of SDRE based methods, and their relative improvement across effective horizon, is attributed to their MSE being polynomial in the effective horizon whereas the MSE of traditional importance sampling techniques based on products of action probability ratios is exponential in the effective horizon. The stronger dependency of IS, PDIS, and INCRIS on the effective horizon is also demonstrated in Figure 3.
Figure 3: Normalised residuals ($y$-axis), defined as $(\hat{G}-\mathcal{G})/H$, as a function of the effective horizon ($H$; $x$-axis) of the taxi domain. Residuals are represented by their mean $\pm$ standard error over 20 independent runs.
Table 4: Mean squared error (MSE) for Monte Carlo estimators of the expected return in the inventory management domain based on 50 independent runs for each effective horizon. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each effective horizon.
Effective horizon
WIS
WSIS
WPDIS
WSPDIS
WINCRIS
WSINCRIS
SDRE
SSDRE
WDR
WDRSIS
10
0.6080
0.4525
0.7234
0.5084
0.4896 0.4767
1.2764
1.2790
53.2003
53.1011
50
15.8227
15.9281
11.6239
10.7777
35.6350
35.8282
9.4938 9.4435
188.8620
187.3876
250
440.6435
473.4986
281.9518
273.3647
1753.7938
1752.6067
25.3142 25.2920
993.3100
1055.8603
1000
8233.7773
7948.8306
4876.1759
4652.7797
27675.6630
27635.4627
231.6101 231.4313
11982.1668
12176.5372
8 Conclusion
Off-policy reinforcement learning comes with the significant challenge of high-variance estimates for the expected return. This paper introduces state-based importance sampling, a class of off-policy evaluation techniques which reduces the variance of importance sampling by decreasing the effective trajectory length through the elimination of “negligible states” for which the chosen action has no impact on the future expected return. Theoretical results on ordinary importance sampling demonstrate how the exponential dependence of the variance and MSE on $H$ can be improved into an exponential dependence on the maximal sub-trajectory length of non-negligible states. We show that either a covariance testing or a Q-value based identification algorithm can be used to identify such negligible states and that IS can be integrated with a variety of off-policy evaluation methods, including per-decision importance sampling, incremental importance sampling, weighted importance sampling, stationary density ratio estimation, and doubly robust estimation. Due to the various base-estimators’ differing assumptions, state-based techniques can be applied to a wide variety of domains by selecting a suitable base-estimator. Experimental results demonstrate the benefits of state-based methods as compared to their traditional counterparts on four distinct domains. Future work directions include theoretical studies on the presence of negligible states, implementing policy optimisation techniques based on state-based importance sampling, and exploring the links with other off-policy evaluation techniques.
Acknowledgements
David M. Bossens was supported by the UKRI Trustworthy Autonomous Systems Hub, EP/V00784X/1. Philip S. Thomas was supported by the National Science Foundation under grant no. CCF-2018372. We also thank Shreyas Chaudhari for comments on a draft of the paper.
Appendix A Appendix A: Proof of $\mathbb{E}[A]=1$
@@
The MSE of SIS is therefore given by
@@
Let $s_{t}\in S^{A}=\{s\in\mathcal{S}:|\hat{Q}_{t^{\prime}}(s,a)-\hat{Q}_{t^{\prime}}(s,a^{\prime})|<\epsilon\forall a,a^{\prime}\in\mathcal{A}\forall t^{\prime}\in[H]\}$. This implies that the SIS estimator has the following expected return from starting state $s_{t}$
@@
where the approximate equality follows since $|\hat{Q}_{t}(s,a)-\hat{Q}_{t}(s,a^{\prime})|<\epsilon$ implies $Q_{t}(s,a)\approx Q_{t}(s,a^{\prime})$ for all $a,a^{\prime}\in\mathcal{A}$. Under a perfect model, the bias of the estimator comes only from the approximate equality, in which case the absolute bias induced by state being dropped is at most $\epsilon$. Under an erroneous model, which yields $|\hat{Q}_{t}(s,a)-Q_{t}(s,a)|<\epsilon_{2}$ for all $t$ and all $a\in\mathcal{A}$, the absolute bias induced by a state being dropped is at most $\epsilon+\epsilon_{2}$. Repeatedly applying this argument for a maximal number of $M_{A}$ dropped weights in a trajectory, the absolute bias induced is at most $M_{A}(\epsilon+\epsilon_{2})$. Filling in $O\left(\text{poly}(H)\right)$ as the bias and $\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{M_{B}}\right)^{2}$ as the variance in the bias-variance decomposition, the MSE is bounded by $O\left(\exp(M_{B})\right)$. ∎
@@
Table 5: Mean squared error (MSE) comparing state-based importance sampling methods with different negligible state sets on 50 independent runs of the deterministic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each number of episodes; in case of a tie only the best estimators are highlighted in bold underlined.
IS
SIS
@@
7
0.0067
0.0019
0.0056
0.0019
9
0.0153
0.0019
0.0138
0.0019
11
0.0353
0.0029
0.0345
0.0029
13
0.0691
0.0026
0.0705
0.0026
15
0.1389
0.0033
0.1386
0.0033
17
0.3026
0.0031
0.3035
0.0031
Table 6: Mean squared error (MSE) comparing state-based importance sampling methods with different negligible state sets on 200 independent runs of the stochastic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size; in case of a tie only the best estimators are highlighted in bold underlined.
WIS
WSIS
@@
7
10.2238
8.4832
10.2186
8.4832
9
9.4737
7.7326
9.4489
7.7326
11
9.7285
6.2044
9.6775
6.2044
13 8.4436
9.5003
8.4058
9.4967
15
18.0453
4.2803
17.9633
4.2831
17
8.2474
4.7482
7.7646
4.8129
Table 7: Mean squared error (MSE) comparing doubly robust methods with different negligible state sets on 200 independent runs of the stochastic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size; in case of a tie only the best estimators are highlighted in bold underlined.
WDR
WDRSIS
@@
7
5.2572
5.0844
5.2662
5.0844
9
2.8067
2.7486
2.8235
2.7486
11
2.3016
1.9838
2.3058
1.9838
13
1.3459
1.3359
1.3612
1.3354
15
1.2258
0.8788
1.1675
0.8797
17
0.8385
0.4817
0.6620
0.4820

### v5

@@ VII-A Lift domains
The deterministic lift domain involves an agent moving on a straight line, with the state space $\mathcal{S}=\{-b,\dots,-1,0,1,\dots,b\}$ representing the coordinates along the line and the action space $\mathcal{A}=\{\texttt{left},\texttt{right}\}$ representing left or right movements on the line. With the exception of the starting state and states within a distance of 1 from the terminal state, the domain has lift states which force the agent to move left on $s<0$ and right on $s>0$. The episode stops when the agent hits the bound, and then the terminal reward of $r=s$ is given (equal to either $-b$ or $b$). All non-terminal rewards are $r=-1$ to penalise the length of the path. The experiments manipulate this bound for an increasing effective horizon. An analysis comparing state-based estimators to their traditional counterparts (see Tab. I) demonstrates a near universal benefit in the deterministic domain, with the only exception being SDRE. SIS is the best performing estimator across all domain sizes, followed by IS and SPDIS depending on the domain size. SDRE and SSDRE have particularly poor estimates because the domain does not have a stationary distribution. Analysis of the residuals (see Fig. 1a in Appendix F) also shows that: a) state-based methods have successfully reduced the variance compared to their traditional counterparts; b) SPDIS strongly reduces the residual of PDIS for large domain sizes; and c) INCRIS and SINCRIS have large residuals, which implies that dropping recent time steps may not be suitable for the deterministic domain (e.g. the first time step always determines the final reward). Due to the deterministic domain being easy to model, the doubly robust methods have MSEs close to zero in this domain so are not further analysed.
Stochastic lift domain
In the stochastic lift domain, the evaluation policy and the state transitions are stochastic rather than deterministic. The evaluation policy $\pi_{e}$ takes the best action with probability $1-\delta$ and the other action with probability $\delta$, and the state transitions dynamics are formulated such that the action effect is reversed with probability $\delta$. An analysis comparing state-based estimators to their traditional counterparts (see Tab. II) demonstrates a benefit in the stochastic domain. This time doubly robust methods are also analysed as the error of the DR estimator is now non-zero. WSIS and WSPDIS consistently outperform their traditional counterparts. For large domain sizes, WDRSIS is the best performing estimator followed by WDR and WSPDIS. State-based methods are not adversely affected by the domain size, which is explained by all the lift states being dropped and therefore the maximal number of steps $M_{B}$ in non-negligible states remaining constant regardless of domain size. For small domain sizes, WINCRIS and WSINCRIS have the best performance, indicating the benefit of dropping time steps; the difference to the deterministic domain maybe due to the visitation of multiple lifts in the same episode. The lower performance of WINCRIS and WSINCRIS in larger domain sizes is attributed to the limited history window $k\in\{1,\dots,10\}$, which is needed for computational feasibility but may not be sufficient for longer trajectories. Fig. 1b in Appendix F further demonstrates that the state-based methods have successfully reduced the variance compared to their traditional counterparts. As in the deterministic lift domain, the poor performance of SDRE estimators in the stochastic lift domain is attributed to the domain not having a defined stationary state distribution.
VII-B Inventory management
The task of the agent in the inventory management problem is to purchase items to make optimal profits selling the items. The agent must balance supply with demand while having at most $S$ items in the inventory. The state is the current inventory while the action is the chosen number of items to add. States corresponding to an almost full inventory are similar to lift states in the sense that the actions chosen are unimportant. As the domain is high variance due to the varying demand, we opt for a relatively high number of 50 experiment runs and instead of manipulating the effective horizon we analyse the improvement in performance across different numbers of episodes. An analysis comparing state-based estimators to their traditional counterparts (see Tab. III) demonstrates a near universal benefit in the inventory management domain, with 2– to 8–fold improvements being observed (see Tab. III) – with the exception of DR which already has high accuracy. WSPDIS performs best for 100 episodes while SSDRE performs best for 250 to 1,000 episodes. The high performance of SSDRE is attributed to a clearly defined stationary distribution. The performance of SSDRE and SDRSIS dependent on the sample size suggests that these state-based methods provide improved convergence and efficiency. Other high-performing algorithms are WSIS, WSINCRIS, WDR, and WDRSIS. Fig. 2 in Appendix F demonstrates that state-based techniques reduce the variance and yield a smaller residual.
VII-C Taxi
The taxi domain has a much larger state space with $|\mathcal{S}|=2000$ and can have infinite horizon. A taxi agent has to pick up and drop off passengers who arrive at random corners in a $5\times 5$ grid based on $6$ actions, namely going a step north, west, south, east, picking up a passenger, and dropping off a passenger. The taxi agent receives a reward of $20$ upon a drop-off or pick-up at the correct location, and a reward of $-1$ otherwise. The domain is of interest to investigate the scalability to large state-action spaces and large effective horizon, and it is of interest to manipulate the effective horizon in an additional domain which in contrast to the lift domains has a stationary distribution. The results for the taxi domain (see Tab. IV) also show a near universal benefit for state-based estimators compared to the traditional counterparts, although the effect is not as pronounced as for example in the inventory management domain. For an effective horizon of 10, WSIS is the highest performing followed by WSINCRIS and WINCRIS. For effective horizons between 50 and 1,000, SSDRE is the highest performer followed by SDRE and WSPDIS. The high performance of SDRE based methods, and their relative improvement across effective horizon, is attributed to their MSE being polynomial in the effective horizon whereas the MSE of traditional importance sampling techniques based on products of action probability ratios is exponential in the effective horizon. As illustrated in Fig. 3 in Appendix F, the SDRE methods have a lower variance than WDR methods and their residuals do not depend as strongly on the effective horizon as WIS, WPDIS, and WINCRIS.
VIII Conclusion
Off-policy reinforcement learning comes with the significant challenge of high-variance estimates for the expected return. This paper introduces state-based importance sampling, a class of off-policy evaluation techniques which reduces the variance of importance sampling by eliminating states that do not affect the return from the importance weight computation. The state-based technique is shown to improve performance when integrated with a variety of off-policy evaluation methods, making it a useful plug-in for many applications. Future work directions include theoretical studies on the presence of negligible states, implementing policy optimisation techniques based on state-based importance sampling, and exploring the links with other off-policy evaluation techniques.
Acknowledgement
We thank Shreyas Chaudhari for comments on a draft of the paper.
Supplementary Information
Appendix A: Proof of $\mathbb{E}[A]=1$
@@ Acknowledgement
With the bias equal to the covariance, it follows via bias-variance decomposition that
@@
Let $s_{t}\in S^{A}=\{s\in\mathcal{S}:|\hat{Q}_{t^{\prime}}(s,a)-\hat{Q}_{t^{\prime}}(s,a^{\prime})|<\epsilon\quad\forall a,a^{\prime}\in\mathcal{A}\quad\forall t^{\prime}\in[H]\}$. This implies that the SIS estimator has the following expected return from starting state $s_{t}$
@@
where the approximate equality follows since $|\hat{Q}_{t}(s,a)-\hat{Q}_{t}(s,a^{\prime})|<\epsilon$ implies $Q_{t}(s,a)\approx Q_{t}(s,a^{\prime})$ for all $a,a^{\prime}\in\mathcal{A}$. Under a perfect model, the bias of the estimator comes only from the approximate equality, in which case the absolute bias induced by state being dropped is at most $\epsilon$. Under an erroneous model, which yields $|\hat{Q}_{t}(s,a)-Q_{t}(s,a)|<\epsilon_{2}$ for all $t$ and all $a\in\mathcal{A}$, the absolute bias induced by a state being dropped is at most $\epsilon+\epsilon_{2}$. Repeatedly applying this argument for a maximal number of $M_{A}$ dropped weights in a trajectory, the absolute bias induced is at most $M_{A}(\epsilon+\epsilon_{2})$. Filling in $\mathcal{O}\left(\text{poly}(H)\right)$ as the bias and $\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{M_{B}}\right)^{2}$ as the variance in the bias-variance decomposition, the MSE is bounded by $\mathcal{O}\left(\exp(M_{B})\right)$. ∎
@@
The comparison for traditional importance sampling includes SIS (Lift states), which uses the lift states of the domain (i.e., where $\mathcal{S}^{A}=\mathcal{S}^{L}$); SIS (Covariance testing), which uses the negligible states identified by covariance testing; and SIS (Q-based), which uses the negligible states identified using Q-values. Similarly, we evaluate the effectiveness of state identification for the doubly robust off-policy evaluation method, which are analgously denoted as SDR (Lift states), SDR (Covariance testing), and SIS (Q-based).
Analysing the different negligible state set identification methods reveals that SIS (Q-based) identifies the lift states as negligible states and therefore yields the same performance as SIS (Lift states). Tab. V demonstrates the superior performance of SIS (Q-based) and SIS (Lift states) compared to IS, which never drops any states, and SIS (Covariance testing), which typically drops two lift states (the maximal cardinality given the computational budget).
TABLE V: Mean squared error (MSE) comparing state-based importance sampling methods with different negligible state sets on 50 independent runs of the deterministic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each number of episodes; in case of a tie only the best estimators are highlighted in bold underlined.
IS SIS
@@
7 0.0067 0.0019 0.0056 0.0019
9 0.0153 0.0019 0.0138 0.0019
11 0.0353 0.0029 0.0345 0.0029
13 0.0691 0.0026 0.0705 0.0026
15 0.1389 0.0033 0.1386 0.0033
17 0.3026 0.0031 0.3035 0.0031
Results for state set identification are analogous to the deterministic lift domain: WSIS (Q-based) is successful at identifying the lift states as negligible states and together with WSIS (Lift states) has a superior performance compared to WSIS (Covariance testing) and WIS (see Tab. VI; and WSIS (Covariance testing) typically drops two lift states (the maximal cardinality given the computational budget). Tab. VII demonstrates how dropping states is also beneficial for doubly robust methods and again the Q-based state identification works to identify the lift states.
TABLE VI: Mean squared error (MSE) comparing state-based importance sampling methods with different negligible state sets on 200 independent runs of the stochastic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size; in case of a tie only the best estimators are highlighted in bold underlined.
WIS WSIS
@@
7 10.2238 8.4832 10.2186 8.4832
9 9.4737 7.7326 9.4489 7.7326
11 9.7285 6.2044 9.6775 6.2044
13 8.4436 9.5003 8.4058 9.4967
15 18.0453 4.2803 17.9633 4.2831
17 8.2474 4.7482 7.7646 4.8129
TABLE VII: Mean squared error (MSE) comparing doubly robust methods with different negligible state sets on 200 independent runs of the stochastic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size; in case of a tie only the best estimators are highlighted in bold underlined.
WDR WDRSIS
@@
7 5.2572 5.0844 5.2662 5.0844
9 2.8067 2.7486 2.8235 2.7486
11 2.3016 1.9838 2.3058 1.9838
13 1.3459 1.3359 1.3612 1.3354
15 1.2258 0.8788 1.1675 0.8797
17 0.8385 0.4817 0.6620 0.4820
Appendix E: Experimental details
In the experiments, covariance testing is applied with $\epsilon_{s}=\epsilon=0.01$ and rather than the full power set, we limit the search to sets with cardinalities of at most 2. The Q-value based identification is suitable for all the domains and the parameter settings are $\epsilon=1$ in the grid world lift domains, $\epsilon=50$ in the inventory management domain, and $\epsilon=2$ in the taxi domain.
In the lift domains, $\pi_{e}$ is the optimal policy and $\pi_{b}$ is the uniform random policy. The expected return is 1.0 in the deterministic domain and negative in the stochastic domain (with magnitude increasing with domain size). The failure probability is set to $\delta=0.05$ for the stochastic domain.
In inventory management, the purchasing cost is set to $2.49$, the sale price to $3.99$, and the holding cost to $0.03$. Demand for items follows a Gaussian with mean $S/4$ and standard deviation $S/6$, and $S=10$ in the experiments. Each episode consists of 100 steps, and consistent with the rest of our experiments an undiscounted setting ($\gamma=1$) is considered. The evaluation policy, $\pi_{e}$, was trained for 500 episodes on the MDP and has an expected return of $\mathcal{G}=1471.1041$. The behaviour policy, $\pi_{b}$, is uniform random.
In taxi, the implementation of the domain is based on code of [7] (see {https://github.com/zt95/infinite-horizon-off-policy-estimation/tree/master/taxi}). The evaluation policy represents a policy formed after 1,000 iterations of Q-learning and the behaviour policy represents a policy formed after 950 iterations of Q-learning. The expected return is positive and increases with effective horizon.
Appendix F: Residual plots
(a) Deterministic lift domain
(b) Stochastic lift domain lift domain
Figure 1: Residuals ($y$-axis), defined as $\hat{G}-\mathcal{G}$, as a function of the domain size ($x$-axis) in the lift domains. Estimates are based on 1,000 episodes. Residuals are represented by their mean $\pm$ standard error over 50 independent runs for the deterministic lift domain and over 200 independent runs for the stochastic domain. Note that SDRE and SSDRE are not included for improved visibility as their residuals are extremely large.
Figure 3: Residuals ($y$-axis), defined as $\hat{G}-\mathcal{G}$, as a function of the effective horizon ($H$; $x$-axis) of the taxi domain. Residuals are represented by their mean $\pm$ standard error over 20 independent runs.